OpenStudy (anonymous):

substitution -2x+y=2 and 2x+y=6

OpenStudy (anonymous):

2x + y = 6 y = 6 - 2x -2x + y = 2 -2x + (6 - 2x) = 2

OpenStudy (anonymous):

continue the process

OpenStudy (anonymous):

how to solve this problem i am new to this lesson?

OpenStudy (anonymous):

-2x + (6 - 2x) = 2 -2x + 6 -2x = 2 combine similar terms, -4x + 6 = 2 transpose 6 to the other side, -4x = 2 - 6 subtract -4x = -4 divide both sides by -4 x = 1

OpenStudy (anonymous):

then just substitute the value of x to any equation you like.

OpenStudy (anonymous):

where did you get the 4?

OpenStudy (anonymous):

auh okey,i understand you multiply it tnx

OpenStudy (anonymous):

Your welcome :D

OpenStudy (anonymous):

did you get the value of y?

OpenStudy (anonymous):

not yet....still getting

OpenStudy (anonymous):

Okay. take your time. Tip: substitute it to where Y is the only variable at one side and all the other at the other side. XD

OpenStudy (anonymous):

is my answer correct? y=2x-6 y=2(1)-6 y=-2-6 y=-8 ????? is it right or wrong i really wanted to learn.

OpenStudy (anonymous):

You're almost right. Just remove the negative sign at 2 . How did you obtain a negative sign at 2?

OpenStudy (anonymous):

ohhh!! no!!!!

OpenStudy (anonymous):

okey sorry.. so the answer would be y=2x-6 y=2(1)-6 y=2-6 y=-8 ???

OpenStudy (anonymous):

when you subtract 2 by 6 you'll get what? it's -4, right? not -8

OpenStudy (anonymous):

yeah.okay sorry about that

OpenStudy (anonymous):

tip: just subtract the lower from the higher then base the sign on the higher number.

OpenStudy (anonymous):

understood. do you know how to do checking??

OpenStudy (anonymous):

yeah. Using the ORIGINAL equation, just substitute both values.

OpenStudy (anonymous):

okey. thank you so much for the tutorial i really enjoy it thanks alot!

OpenStudy (anonymous):

No problem. if you need help, don't hesitate to ask. :]]

OpenStudy (anonymous):

ok.:)